2 divided by 2/X + 2/Y is problem one...i have an answer guide so i know the answer i just don't know how to do it on my own
2x^2/3y * y^3/8x is problem two... this too i have the answer to but i don't know how to do it myself

Answer by jainenderkapoor(61)

Q.1 2 divided by (2/x + 2/y)

(2/x + 2/y) = 2 (1/x + 1/y)
= 2 ( (y + x)/xy)
(I solved 1/x + 1/y by finding common denominator xy and then multiplying the numerator by x and y respectively)
= 2(y+x)/xy
Now we have to solve 2 divided by 2(y + x)/xy
we get 2 xy/2(y + x)
= xy/(y + x)

Q.2 (2x^2/3y)* (y^3/8x)

Here we will simply multiply the numerators and the denominators
(2x^2 * y^3)/(3y*8x)
we get the power of x as 2 - 1 = 1 in the numerator and the power
of y as 3 - 1 = 2 in the numerator
We get 2/24 which can be written as 1/12
so the final answer becomes
(x y^2)/12

I hope that the solution is clear to you.
If you have any doubt you are welcome to contact me.
I am also available for online tutoring in math at very reasonable rates.
My mail id is jainenderkapoor@gmail.com
Jainender Kapoor